I got the equation to use for unemployment rate but I need help with the rest.

In alphaland there are 275,000 employed; 25,000 were unemployed; and 100,000 were not in the labor force. Assume all citizens are adults. Use this information to calculate:
a. Adult population

b. Labor force

c. Labor force participation rate

d. Unemployment rate 6.25%
Answer #1

Answer
a. Adult population
Adult population =employed +unemployed +not in labor force
=275000+25000+100000
=400000
b. Labor force
Labor force =employed +unemployed
=275000+25000=300000

c. Labor force participation rate
Labor force participation rate =(Labor force/ Adult population)*100
=(300000/400000)*100
=75%

d. Unemployment rate

Unemployment rate =(unemployed/labor force)*100
=(25000/300000)*100
=8.33333333
=8.33%

======
The unemployment rate is unemployed of the labor force and not population so the unemployment rate is 8.33%